# Vid2Fluid: 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting

### v1

@@ 3.3.1. Preprocess with Generative Gaussian Splatting
information, we rely on the existing single-image to 3DGS generation methods (Zou et al., 2024; Xiang et al., 2024). However, these methods are trained on datasets of 3D surface representation and the generated 3D Gaussians are typically sparse and concentrated on the object’s surfaces. The generated 3DGS need to be denoised and densified so that the Gaussian points could precisely carry every pixel information from the frame. As shown in the inset, we sample the 3DGS space into grids and insert a Gaussian at the center of the grid $p_{c}$ if $p_{c}$ is inside the fluid by checking if over half of the number of the intersections of 3DGS’s outer hull and the rays from $p_{c}$ is odd. Additionally, the Gaussian primitives with low opacity or heavily stretchered covariance are pruned.
@@
We also need to carefully deal with the consistency of geometry for the generated 3D Gaussians in continuous frames. Generative methods do not guarantee that the output variation is on the same scale with the perturbation of input conditions. Moreover, adjacent frames could be generated with quite different geometries at the backside from the camera, though the foreground is conditioned by similar inputs. This is acceptable for free-moving fluid like smoke, which shows highly dynamic behavior in the video. However, this inconsistency can impact the calculation of physically grounded fluid dynamics for flowing rivers on fixed riverbeds. To address the issue, we perform the union operation on the generated 3D Gaussians from $N$ consecutive frames, forming a batch to be used in the subsequent process of simulation parameter optimization in Sec. 3.4. Note that $N$, the number of frames to union, is dynamically decided by the moving intensity of the object profile. This is achieved by evaluating similarities between adjacent frames $f$ and $f+1$ throughout the video: $N\propto MSE\big{(}PSNR({f,f+1})\big{)}$.
3.3.2. 3D Fluid Surface Velocity Estimation
@@
After obtaining the geometry, we can estimate the volumetric velocity field of the object. This starts by first estimating the 3D surface velocity, where we utilize optical flow (Farnebäck, 2003) in the screen space, as pixels moving on the screen can be seen as material points being displaced in the Normalized Device Coordinates (NDC). Additionally, we combine the $z$-depth information from the preprocessed 3DGS to get displacements in the third dimension. However, optical flow heavily relies on the color change of pixels between frames, which is not detectable in the area that exhibits a uniform flow, where no splashing or vortex effects could be exploited. To tackle this problem, we propose a mainstream direction detection method and compensate the region where optical flow fails. The mainstream direction is determined by edge orientations of fluid detected from the video, and then mapped to the entire region of fluid. When filling in the missing parts of the optical flow result, we project neighboring optical flows in the mainstream direction, with those more aligned with the main direction receiving greater weighting.
(5) $\vec{d}_{opz,k}=\sum_{i\in B(k)}w(i)\max(0,\bigg{(}\vec{n}_{k}\cdot\frac{\vec{d}_{opz,i}}{|d_{opz,i}|}\bigg{)})\cdot\vec{d}_{opz,i},$
@@
where $\vec{d}_{opz}$ is the 3D velocity in NDC space, calculated from optical flow and depth change. $i$ denotes the detectable place by optical flow, and $k$ is the missing part. $B$ bounds the range of $i$ contributing to specific $k$ and defines certain distance-based weight $\omega(i)$. The dot product of mainstream direction $\vec{n}_{k}$ and the normalized $\vec{d}_{i}$ measures the cosine value of their angle.
@@
We then project the velocity on the screen space to the 3D Gaussian space: $\vec{v_{p}}=\big{(}S\circ(PW)_{+}T\big{)}^{-1}\circ\vec{d}_{opz,i}$ , where $\vec{v}_{p}$ is the particle velocity, $P$ is the view projection matrix, and $W$ is the world-to-camera matrix. $T$ and $S$ denote the transition and scaling same as those (if any) on input images before generating the 3DGS.
$3$D Volumetric Velocity Estimation
Figure 3. Illustration (left) and a real volumetric velocity field (right). The velocity is dampen near the riverbed. On the right figure, the darker the color, the smaller the velocity.
With the surface velocity extracted, we interpolate it across the entire volume of preprocessed Gaussian particles to initialize the dynamic behavior of the fluid object. Before interpolation, the enclosure of the fluid must be determined, especially in the regions of 3D Gaussians that are unseen from the camera. Depending on the constraint strength of the boundary, we let the surface velocity determine relative magnitudes and impose the presumed velocity profile. For flowing liquid, the boundary layer approximation is imposed on the riverbed as shown in Fig. 3, but for freely moving fluid, we do not dampen the velocity on the unseen sides. This ensures a gradual change in the velocity of the flowing layer as it approaches any solid boundary, where the velocity should ultimately be zero. To reveal this effect, we choose the third-order polynomial of the integral approximate solution of laminar external flows:
(6) $\frac{|\vec{v}|}{V_{\infty}}=\frac{3}{2}\left(\frac{y}{\delta}\right)-\frac{1}{2}\left(\frac{y}{\delta}\right)^{3},$
@@
where $V_{\infty}$ is the magnitude of corresponding surface velocity, $\delta$ is the boundary layer thickness, which is dynamically decided by the fluid type: for a flowing river it is set to a larger thickness, while for smoke, it is limited to a sub-cell scale so it introduces very small damping. $\vec{v}$ takes the direction of the mainstream but not the same as its surface counterpart. After enclosing the surface of the object, we calculate the inner velocity field by interpolation weighted as a Gaussian distance distribution and finalize it by another divergence-free projection in the next step.
3.4. Presumed APIC Simulation and Optimization
3.4.1. Grid-Based Velocity Evolution
@@
We transfer the velocity field from particles to the grid, which offers two main advantages. First, grid-based representation helps us more efficiently track the fluid movement between two frames. Second, by solving Eq. 3 alongside grid-based convection, particle velocity can be transferred to grid using Eq. 2 with $\mathbf{C}$ set to zero, making the dynamics of Gaussians differentiable.
To formulate the optimization problem , we modify two parts of Eq. 3. First, to solve the Poisson Pressure Equation, instead of using iterative methods such as Jacobi iterations, we implement and modify the PeRCNN (Rao et al., 2021) structure shown in the inset where $P^{(k)}$, $B.C.$ and $\vec{v}$ are the latents of pressure , boundary condition and velocity respectively at $k$-th recurrent layer.
@@
The approach is as accurate as the numerical methods while facilitates the convergence of the optimization. Secondly, as the original APIC simulation method does not solve the convection part on the grid, we additionally integrate the back-tracing mechanics from stable flow (Stam, 1999) in the forward process, allowing for unconditional stability in resolving grid-based velocity. However, since back-tracing involves grid indexing and round-off operations, which reduce the differentiability of the optimizing process, we directly use the analytic gradient of convection term of the split PDE: $\frac{\partial\vec{v}}{\partial t}=-\vec{v}\nabla\vec{v},~{}\frac{\partial\vec{v}}{\partial v_{i}^{{}^{\prime}}}=-dt\frac{\partial\vec{v}}{\partial i},$ where $\vec{v}$ denotes the grid velocity after convection, $v_{i}^{{}^{\prime}}$ denotes the velocity component of dimension $i$ before convection.
3.4.2. Loss Design and Parameter Optimization
@@
(7) $\vec{v}_{sim}^{i+n}=\text{DiffSim}^{n}$ $\big{(}\vec{v}^{i}_{gt};v_{in},\tilde{v}_{in},v_{out},\rho,\nu,(b,d)|_{{}_{B.C.}},g,dt\big{)}$
$\min~{}\text{gridMask}\circ$ $\bigg{(}A(\frac{\vec{v}_{sim}}{|v_{sim}|}\cdot\frac{\vec{v}_{gt}}{|v_{gt}|})+BL_{2}(\vec{v}_{sim},\vec{v}_{gt})\bigg{)}.$
3.4.3. Augmentation of Optimized Result
@@
Here we approximate the affine matrix $C$ as the velocity gradient and update the deformation gradient $F$ each time step, which then deforms 3D Gaussians and is reflected by the change of its covariance $A$.
4. Results
We implement the framework of Vid2Fluid mainly using Python on a workstation PC equipped with a 24-core Intel(R) Xeon(R) Platinum 8255C CPU and an NVIDIA Tesla V100 GPU. For the 3DGS generation part, we port the open-source implementation of the existing methods (we use TriplaneGaussian (Zou et al., 2024) by default), and output intermediate states. For re-simulating the flowing fluid, we implemented a GPU-parallelized APIC method with Taichi (Hu et al., 2019). The rendering of obtained 3D Gaussians is compatible with a general 3DGS visualizer. We also implement a relighting interface using GaussianShader (Jiang et al., 2023) to achieve higher resolution and incorporate transparent fluid rendering within the original 3DGS CUDA rasterizer.
4.1. Ablation Studies
4.1.1. Denoising and Filling of 3D Gaussians
For the image-conditioned 3DGS, downsampling (e.g., reducing the image resolution to $256\times 256$ pixels in TriplaneGaussian) results in sparse Gaussian points, each with a large elliptical shape. This is undesirable when relating continuous surface velocities with such sparse 3D Gaussians. Thus, as shown in Fig.4, we preprocess the generated Gaussians, which leads to a more compact and accurate representation as mentioned in Sec. 3.3.
Figure 4. Proprecessing operations. The denoising operation removes outlier Gaussian points, returning a smoother geometry (left bottom) compared the geometry on left top. The filling operation inserts Gaussians into sparse generated 3DGS (middle top) and output a dense 3DGS (middle bottom). Our union strategy merges the generated 3DGS from multiple frames and outputs a higher-fidelity geometry (right bottom) compared direct union (right top).
4.1.2. Mainstream-corrected Optical Flow
As mentioned in Sec. 3.3.2, optical flow may return a 2D velocity field with “vacant” static regions. Fig. 5 shows that the mainstream-corrected result accurately captures the true motion where the original optical flow detects near-zero velocities.
Figure 5. The mainstream guided optical flow result (right) presents a more meaningful velocity field compared the unguided result where the velocity is vanished in regions such as the one in the closeup (left).
4.1.3. Neural Surrogate for Poisson Pressure Equation
Fig. 6 demonstrate that our modified PeRCNN (Rao et al., 2021) architecture for solving the Poisson pressure equation is equivalently accurate compared with traditional numerical iterative method. Also, the comparison between the optimization processes of PeRCNN and Jacobi iterations reveals that PeRCNN achieves faster convergence.
Figure 6. Left: We demonstrate that our PeRCNN-based optimization converges in the case of honey. Right: We compared our method with the traditional numerical method to show that our method achieves faster convergence.
4.2. Evaluation on Different Fluid Types
We apply our method to three different types of fluids – gas, inviscid liquid, and viscous fluid – to show the generalization of our framework. For the gas case, we compared our method with HyFluid (Yu et al., 2023) in terms of reconstruction quality and novel view synthesis. It is important to note that HyFluid requires multi-view images as input, whereas our approach relies on only single-view images. As shown in Fig. 7, our result is visually more similar to the input image, thanks to the high-degree harmonic features of 3DGS that align effectively with the input image. In terms of novel views, our framework incorporates a generative process, while HyFluid directly takes as input images from different viewpoints, which allows it to achieve higher similarity to the true smoke in these cases.
Figure 7. Comparison with Hyfluid. Top: Input video frames. Middle: Hyfluid output. Botton: Our output dynamic fluid.
The effects of gravity and boundary conditions are more pronounced in fluids with well-defined shapes. For viscous fluids, the relative motion between different parts is constrained by viscosity and further dampened by the boundary. As a result, by using the same governing equations with different scales of optimized simulation parameters, the reconstructed fluid can effectively mimic a wide range of dynamics observed in the video as shown in Fig. 10.
4.3. Video Velocity Reconstruction
Two synthetic videos from known fluid data are used to compare the reconstructed 2D velocity fields with ground truth. As Vid2Fluid inputs planar velocity that is discretized by frame rate and scaled to screen dimensions, both ground truth and reconstructed 3D velocities are projected into 2D NDC space for comparison for visual examination. Figure 8 shows results for two fluid types. The reconstructed in-plane velocity can capture meso-scale profiles but retains noises on fluid surfaces. Optimizing physics-based simulation and aligning velocity direction yields smoother results and enhances physical details, such as boundary collisions in the river and strain development in honey inflow.
Figure 8. Comparison of planar velocity for fluid as river (top) and honey (bottom). Left column is the ground truth, middle column is the result by Vid2Fluid, and right column is the estimated planar velocity in early stage of the framework as described Fig.2.
4.4. Performance with Different 3DGS Generation Methods
As a generic framework based on generative 3DGS, Vid2Fluid is not dependent on a specific method of Gaussian generation. We test our method with different 3DGS generation methods: generative methods, TriplaneGaussion (TriGS) and TRELLIS (Xiang et al., 2024), and multi-view reconstruction method, vanilla 3DGS (Kerbl et al., 2023). These methods differ in terms of image fidelity, resolution, and creativity, thus catering to varying user needs in real-world applications. As shown in Fig. 11, multi-view optimized 3DGS exhibited the highest quality and fidelity with the input images, but its usage in real-world scenarios is limited. TriGS benefits from triplane encoding of image textures, which aligns well with the input image, though this comes at the cost of lower resolution. TRELLIS shows more divergence from the input images, even at high Classifier Free Guidance (CFG) scales, but produces Gaussians at a higher resolution compared to TriGS.
4.5. Different Rendering Options

### v2

@@ 3.3.1 Preprocess with Generative Gaussian Splatting
Since our input only contains single-view information, to obtain the 3D information, we rely on the existing single-image to 3DGS generation methods [ZYG24, XLX24]. However, these methods are trained on datasets of 3D surface representation and the generated 3D Gaussians are typically sparse and concentrated on the object’s surfaces. The generated 3DGS need to be processed so that the Gaussian points could precisely carry every pixel information from the frame.
The directly generated 3DGS exhibits smooth appearance variations across different viewpoints but suffers from coarseness due to input image compression as shown in the inset left. Large elliptical shapes are thereby produced, which is undesirable when relating continuous surface velocities with such sparse 3D Gaussians. To address this, we perform a fast single-view optimization using frames from the input video to enhance the resolution of the reconstructed 3DGS (Fig. 2) While using only one viewpoint significantly reduces computational time, it necessitates a carefully designed loss function to mitigate overfitting due to limited supervision:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{img}+\lambda_{2}\mathcal{L}_{aniso}+\lambda_{3}\mathcal{L}_{vol}+\lambda_{4}\mathcal{L}_{scl}+\lambda_{5}\mathcal{L}_{lumi}.$ (4)
$\mathcal{L}_{img}$ denotes the image loss adopted from original 3D Gaussian Splatting [KKLD23]. $\mathcal{L}_{aniso}$ (anisotropic loss) and $\mathcal{L}_{vol}$ (volume loss ) are implemented the same as Gaussian SPlashing [FFS25]. $\mathcal{L}_{scl}$ represents the geometric mean of Gaussian ellipsoids’ three scale components. This term encourages finer particle details during optimization. $\mathcal{L}_{lumi}$ is the luminance loss, penalizing oversaturated colors in harmonic features. During optimization, we prune Gaussians whose colors closely match the background. This strategy effectively eliminates the misrepresentation of rocks or reefs near the water surface. Meanwhile, this optimization uses a lower threshold than the original 3DGS to encourage densification.
Figure 2: Preprocessed results for generated 3DGS. The left column is the direct generation, and the right column is processed with the single-view optimization.
Subsequently, the optimized 3DGS are denoised and densified. The Gaussian primitives with low opacity or heavily stretched covariance are first pruned. As shown in the inset, we then sample the 3DGS space into grids and insert a Gaussian at the center of the grid $p_{c}$ if $p_{c}$ is inside the fluid by checking if over half of the number of intersections of 3DGS’s outer hull and random rays from $p_{c}$ is odd.
@@
We also need to guarantee the consistency of geometry for the generated 3D Gaussians in continuous frames. Generative methods do not guarantee that the output geometry varies as continuously as input images. Moreover, adjacent frames could be generated with quite different geometries at the backside from the camera, though the foreground is conditioned by similar inputs. This is acceptable for free-moving fluid like smoke, which shows highly dynamic behavior in the video. However, this inconsistency can impact the calculation of physically grounded fluid dynamics for flowing rivers on fixed riverbeds. To address the issue, we perform the union operation on the generated 3D Gaussians from $N$ consecutive frames, forming a batch to be used in the subsequent process of simulation parameter optimization in Sec. 3.4. The frame number $N$ is determined dynamically based on the motion intensity of fluid objects in the video. This is achieved by evaluating similarities between adjacent frames $f$ and $f+1$: $N\propto MSE\big(PSNR({f,f+1})\big)$. The effects of filling and union are seen in Fig. 3. Filling aims to fill in the inner vacancy of the generated fluid body that prevents unreal collapse in the simulation. Union mainly aims to make fluid geometry consistent across frames, especially for the riverbed, which is supposed to be invariant during the flow motion.
Figure 3: The filling operation inserts Gaussians into sparse generated 3DGS in (a) and output a dense 3DGS in (b). Our union strategy merges the generated 3DGS from multiple frames and outputs a higher-fidelity geometry in (c).
3.3.2 Fluid Surface Velocity Estimation
@@
After obtaining the geometry, we can estimate the volumetric velocity field $V_{vol}$ of the object. This starts by first estimating the 3D fluid surface velocity $V_{surf}$, where we first utilize optical flow [Far03] to detect the 2D velocity field in the screen space, denoted as $V_{2D}$, as pixels moving on the screen can be seen as material points being displaced in the Normalized Device Coordinates (NDC). Additionally, we combine the $z$-depth information from the preprocessed 3DGS to get displacements in the third dimension. However, optical flow fails to detect the velocity $V_{2D}$ in locally homogeneous fluid textures, often yielding zero velocity. The detected pixel motion often corresponds to waves or splashing – waves partially reveal the main flow direction, while splashing provides only local information.
To tackle this problem, we propose a mainstream correction strategy and compensate the region where optical flow fails. The process comprises two stages: (1) mainstream-guided neighboring interpolation and (2) physics-constrained velocity correction. Stage (1): In regions where optical flow detection fails, the underlying motion typically exhibits smoothness and reflects bulk fluid behavior. We determine the mainstream direction using geometric constraints (e.g., river bank alignment) and estimate its magnitude from neighboring optical flow, which ought to have comparable kinetic energy.
$\vec{v}_{opz,k}=\sum_{i\in B(k)}w(i)\max(0,\bigg(\vec{n}_{k}\cdot\frac{\vec{v}_{opz,i}}{|\vec{v}_{opz,i}|}\bigg))\cdot\vec{v}_{opz,i}.$ (5)
@@
$\vec{v}_{opz}$ is the 3D velocity in NDC space, calculated from optical flow and depth change. $\vec{n}$ is the mainstream direction, $i$ denotes the detectable place by optical flow, and $k$ is the missing part. $B$ bounds the range of $i$ contributing to specific $k$ and defines a distance-based weight $\omega(i)$. The dot product of mainstream direction $\vec{n}_{k}$ and the normalized $\vec{d}_{i}$ measures the cosine value of their angle.
Unlike elastic bodies, interpolation alone is unsuitable for the amorphous features of liquids. In Stage (2), we derive constraints, based on the divergence-free property of the 3D velocity $V_{surf}$, that the 2D screen velocity $V_{2D}$ must satisfy:
$\nabla_{2D}\cdot\vec{v}_{2D}=-\frac{1}{z}(u\frac{\partial}{\partial u}+v\frac{\partial}{\partial v}+2)v_{z},$ (6)
where $u,v$ are screen coordinates, and $z$ is the depth. The detailed derivation can be found in the Appendix (Section 1.1). This constraint depends exclusively on the out-of-plane velocity $v_{z}$.
In regions where optical flow fails to detect the motion, this constraint cannot uniquely determine the 2D velocity $V_{2D}$ because the optical flow data does not form a closed and well-posed boundary. To resolve this, we employ the projection method with $V_{2D}$ initialized in the aforementioned mainstream direction. We apply this mainstream correction only to areas where optical flow fails, thereby preserving existing turbulence. Based on this interpolation, we then project $V_{2D}$ to satisfy the constraint in Eq.6. Fig. 4 shows that the mainstream-corrected result effectively captures the true motion where the original optical flow detects near-zero velocities, and interpolation alone leads to noisier dynamics.
@@
Finally, the 2D velo city $V_{2D}$ can be inversely mapped to the 3D Gaussian space according to the camera parameters $\vec{v_{surf}}=\big(S\circ(PW)+T\big)^{-1}\circ\vec{v}_{2D}$ , where $\vec{v}_{surf}$ is the 3D velocity of surface particles, $P$ is the view projection matrix, and $W$ is the world-to-camera matrix. $T$ and $S$ denote the transition and scaling same as the preprocess (if any) done on input images before generating the 3DGS. This approach yields physically plausible surface velocities $V_{surf}$ for subprocesses.
Figure 4: The unguided result shows vanished velocity in regions such as the one highlighted in the close-up (left). After applying mainstream-guided neighboring interpolation, we obtain the result shown in the middle. With further physics-constrained velocity correction, we achieve a more meaningful velocity field (right).
3.3.3 3D Volumetric Velocity Estimation
With the surface velocity $V_{surf}$ extracted, we calculate the entire volumetric velocity $V_{vol}$ of preprocessed Gaussian particles to initialize the dynamic behavior of the fluid object. Velocity on the enclosure surface, $V_{encl}=V_{surf}\cup V_{base}$, needs to be fully determined before computing $V_{vol}$. The surface velocity $V_{surf}$ from the previous stage is used, with additional unknowns $V_{base}$ at the riverbed. While the analytical solution requires zero velocity near solid boundaries, discretization allows for damping, meaning that the velocity at the grid cells closest to the boundary does not vanish in the simulation. We employ wall functions from computational fluid dynamics (CFD) to estimate this damping, which depends on grid resolution and fluid type. Although this assumes external laminar flow—limiting fine details—it avoids introducing nonphysical supervision in later optimization. This ensures a gradual change in $V_{base}$ of the flowing layer as it approaches any solid boundary, as shown in Fig. 5. To reveal this effect, we choose the third-order polynomial of the integral approximate solution of laminar external flows:
$\frac{|\vec{v}_{base}|}{V_{\infty}}=\frac{3}{2}\left(\frac{y}{\delta}\right)-\frac{1}{2}\left(\frac{y}{\delta}\right)^{3},$ (7)
@@
where $V_{\infty}$ is the magnitude of corresponding surface velocity, $\delta$ is the boundary layer thickness, which is dynamically decided by the fluid type: for a flowing river it is set to a larger thickness, while for smoke, it is limited to a sub-cell scale so it introduces very small damping. $\vec{v}_{base}$ takes the direction of the mainstream but not the same as its surface counterpart.
Figure 5: Illustration (left) and a real volumetric velocity field (right). The velocity is dampen near the riverbed. On the right figure, the darker the color, the smaller the velocity.
After obtaining the enclosure surface velocity $V_{surf}$, we use a constrained projection method (Eq. 8) to solve for a divergence-free 3D velocity field, with the aforementioned boundary correction providing a closer initialization.
$C_{i}:~(\nabla\cdot\vec{v})_{i}=0,~~~~~~~~~\vec{v}_{j}^{n+1}=\vec{v}_{j}^{n}+\Delta\vec{v}_{j},$ (8)
$\Delta\vec{v}_{j}=\Sigma_{i}\lambda_{i}\nabla_{v_{j}}C_{i},~~~~~~~\lambda_{i}=-\frac{C_{i}(\vec{v}_{1},\dots,\vec{v}_{n})}{\Sigma_{k}|\nabla_{v_{k}}C_{i}|^{2}+\epsilon}.$
For each grid cell $i$, the divergence-free constraint $C_{i}$ depends on its neighboring velocities $\vec{v}_{k}$. During each iteration of the constrained projection solver, the velocity update $\Delta\vec{v}_{j}$ is computed based on the gradients $\nabla_{v_{j}}C_{i}$ of all constraints influenced by $\vec{v}_{j}$. The relaxation parameter $\epsilon$ ensures stable convergence.
3.4 Presumed APIC Simulation and Optimization
3.4.1 Grid-Based Velocity Evolution
@@
We transfer the velocity field from particles to the grid, which offers two main advantages. First, grid-based representation helps us more efficiently track the fluid movement between two frames. Second, by solving Eq. 2 alongside grid-based convection, particle velocity can be transferred to grid using Eq. 1 with $\mathbf{C}$ set to zero, making the dynamics of Gaussians differentiable (see Appendix, Section 2.1 for details).
To formulate the optimization problem , we modify two parts of Eq. 2. First, to solve the Poisson Pressure Equation, instead of using iterative methods, we implement and modify the PeRCNN [RSL21] structure shown in the inset where $P^{(k)}$, $B.C.$ and $\vec{v}$ are the latents of pressure, boundary condition and velocity respectively at $k$-th recurrent layer. PeRCNN replaces the implicit pressure solver during optimization. As demonstrated by PDE-Net [LLMD18], a convolutional kernel with properly designed weights can effectively discretize spatial operations equivalent to those in numerical methods. Different weight configurations can approximate various differential operators (e.g., gradient, divergence, Laplacian), and larger stencil sizes typically enable higher-order discretization schemes. Beyond the convolutional structure, an outer recurrent layer is formed to emulate the iterative process of implicit solvers. This leverages the RCNN structure to facilitate optimization while maintaining effective pressure correction, without requiring an ad-hoc PeRCNN model.
We employ this architecture to solve the Poisson pressure equation in our physical optimization framework, leveraging its differentiable properties and improved stability for convergence. Prior to implementation, we first validate the structure’s numerical accuracy and computational efficacy. This is done by sampling data from our reconstructed 3D velocity field with initialized physical properties, recording the pressure field both before and after solving via an implicit solver. These paired data samples are then used to evaluate the PeRCNN, which is expected to predict pressure results the same as the second-order Laplacian discretization scheme used in the numerical method. As shown in Fig. 6, the validation demonstrates successful convergence; we consequently employ the fitted kernel to solve pressure in subsequent physical optimization, whose efficacy has been demonstrated in the paper.
Figure 6: Optimization network (left) and parameter fitting of its convolutional kernel for pressure solving.
@@
Second, as the original APIC simulation method does not solve the convection part on the grid, we additionally integrate the back-tracing mechanics from stable flow [Sta99] in the forward process, allowing for unconditional stability in resolving grid-based velocity. However, since back-tracing involves grid indexing and round-off operations, which reduce the differentiability of the optimization process, we directly use the analytic gradient of the convection term of the split PDE: $\frac{\partial\vec{v}}{\partial t}=-\vec{v}\nabla\vec{v},~\frac{\partial\vec{v}}{\partial v_{i}^{\prime}}=-dt\frac{\partial\vec{v}}{\partial i},$ where $\vec{v}$ denotes the grid velocity after convection, $v_{i}^{\prime}$ denotes the velocity component of dimension $i$ before convection. For first-order spatial discretization, we employ the upwind scheme to enhance stability.
Fig. 7 demonstrates that our modified PeRCNN [RSL21] architecture for solving the Poisson pressure equation is equivalently accurate compared with traditional numerical methods using iterative solvers or analytical gradients. Also, the comparison between the optimization processes of PeRCNN and Jacobi iterations reveals that PeRCNN achieves faster convergence.
Figure 7: We compared the optimization process whose Poisson pressure equation is solved by PerCNN structure with those solved by (left) Jacobi iteration and (right) analytical gradient. The latter two methods apt for different batch sizes. Both comparisons demonstrate that utilizing the PeRCNN structure leads to faster convergence.
3.4.2 Loss Design and Parameter Optimization
@@
$\vec{v}_{sim}^{i+n}=\text{DiffSim}^{n}$ $\big(\vec{v}^{i}_{gt};v_{in},\tilde{v}_{in},v_{out},\rho,\nu,(b,d)|_{{}_{B.C.}},g,dt\big)$ (9)
$\min~\text{gridMask}\circ$ $\bigg(\alpha(1-\frac{\vec{v}_{sim}}{|\vec{v}_{sim}|}\cdot\frac{\vec{v}_{gt}}{|\vec{v}_{gt}|})+\beta L_{2}(\vec{v}_{sim},\vec{v}_{gt})\bigg).$
During the experiment, each training sample could require up to 25 continuous steps, with about 100 recurrent sub-steps per simulation step, creating a deep and lengthy computation graph. Gradient vanishing would occur in cases like zero initialization, convection schemes requiring indexing, even longer optimization steps, or large surface masks. To address this, optimization is carefully designed with reasonable initialization (constant in/out velocity, non-zero gravity, unity relaxation, CFL-satisfying time step - applicable to multiple fluids with similar scales) and upwind gradient surrogates in convection schemes. Step lengths are also adjusted based on scene dynamics (longer for steady-state, shorter for highly dynamic fluids), and results improve with a higher surface/volume ratio in prior 3D velocity recovery. Physical parameters must be normalized during training. Each parameter is rescaled through an activation function (either exponential or sigmoid, depending on whether negative values are permissible) and then mapped to its corresponding physical scale. This normalization enables more balanced learning rates across parameters.
3.4.3 Augmentation of Optimized Result
@@
Here we approximate the affine matrix $C$ as the velocity gradient and update the deformation gradient $F$ each time step, which then deforms 3D Gaussians and is reflected by the change of its covariance $A$. An example of the re-simulation of our final output geometry and physics is shown in Fig. 8.
Figure 8: The middle figure shows the re-simulation of our final output geometry and physics based on the input video on the left, which is then rendered in the figure on the right.
3.5 Result Rendering